What is the maximum kinetic energy of an ejected electron if platinum metal is irradiated with 166-nm light? The threshold wavelength for a platinum metal surface is 212 nm. Give your answer accurate to three significant figures. Use exponential notation (e.g. 1.23e-19).
Answer #1

threshold energy = h*c/lambda0
= (6.626*10^-34 * 3*10^8)/(212*10^-9)
=9.3764*10^-19 J

Irradiated energy = h*c/lambda
= (6.626*10^-34 * 3*10^8)/(166*10^-9)
= 1.197*10^-18 J

maximim KE = 1.197*10^-18 J - 9.3764*10^-19 J
= 2.59*10^-19 J
Answer: 2.59*10^-19 J